Tuesday 12 July 2011

HBA installation progress — 12-Jul-2011

Today was a slow day on the tile front. Although we got off to a prompt start, we did not receive any new deliveries today and the tile stock is now depleted. We must wait for another shipment before work can be resumed there.

However, work on framesets has continued apace. We have put as many onto the field as we can. There is a limit to this, as we cannot restrict subsequent tile placement. So the excess frames are being stockpiled to one side and will be put on the field as soon as the new tiles go in. Currently the situation looks like this:
We reached this situation by late morning, so then the work effort turned to other things. Some effort was made to correct a rough patch of ground at position H60, and a more general effort was made to smooth out any rubble that was raised by drilling works earlier.

We have also now prepared the ground at the site of the RF container (approx. 10 metres north of antenna H22). The ground has been levelled there and backfilled with a smooth soil level, ready to accept the concrete foundation blocks in the morning.

We are also starting to distribute the posts for the reindeer fence that will encircle the entire site (both HBA and LBA).

One little surprise today was a couple of visitors from the University of Lancaster. Thanks for calling in to say hello!

This evening, we are going to arrange cable ducting for the cable mausoleum and draft up the plans for the RF-container/mausoleum placement. Oh... and update the weblog too.

No comments:

Post a Comment

Note: only a member of this blog may post a comment.